Raffles Bank pays 4% simple interest on its deposit account, whereas Bugis Bank pays interest on its deposit account compounded monthly.
Formulate the quoted and effective interest rates that Bugis Bank should set if it wants to match Raffles Bank, assuming a 5-years horizon period
1
Expert's answer
2021-10-06T09:42:53-0400

Solution:

Interest rate offered by Raffles bank = 4% p.a. simple rate

Time horizon = 5 years

Effective rate of 5 years = simple interest rate "\times" number of years = 0.04 "\times" 5 = 0.20 i.e. 20%

If bugis bank wants to set the interest rate that matches Raffles bank the quoted or nominal rate to set will be the 4% simple interest annually.


To derive the effective interest rate, we use the below formula:

Effective interest rate = "(1 + \frac{i}{n})^{n} -1"


Where: i = Nominal or quoted interest rate = 4%

           n = No. of compounded periods = 12


Effective interest rate = "(1 + \frac{0.04}{12})^{12} -1 = 4.07%"